OpenStudy (blackops2_143):

(a) William measures the dining room ceiling to see how much molding he will need to go around the perimeter of the room’s ceiling. William measured the length of the dining room to be ft and the width to be ft. How much molding does William need for the dining room ceiling? (b) William’s wife measures the window and adds the measurement to the dining room ceiling measurement. She tells her husband that they need a total of ft of molding to remold the ceiling and the window. How much molding is needed for the window?

OpenStudy (blackops2_143):

(c) William begins to plan his work on the window and finds that the pieces for the height of the window will measure ft each. How long will he cut each piece of molding for the width of the window? (d) William’s wife says the molding for the window will take almost 50% of the total molding they need. William says it is closer to of the molding. Approximately what percent of the total molding will be used for the window and who is closer to the correct answer?

OpenStudy (blackops2_143):

i have A,B and C figured out

OpenStudy (blackops2_143):

A: 2(8.75) + 2(10.875) = 39.1/4 feet B: 65 1/8 - 39 1/4 =65 1/8 - 39 2/8 = 25 7/8 feet C: 25 7/8 - (6 3/4 * 2) = 12 3/8 feet

OpenStudy (xofallenangelox):

I have not learned much about these types of equations, so I can't help. I'm sorry D:

OpenStudy (blackops2_143):

oh ok

OpenStudy (blackops2_143):

thx anyways